# REVS: Unlearning Sensitive Information in Language Models via Rank Editing in the Vocabulary Space

### v5

@@ 6 Related Work
Privacy concerns have given way to development of methods preventing regurgitation of PII. Differential privacy (DP; Dwork et al., 2014; Ramaswamy et al., 2020; Hoory et al., 2021; Li et al., 2022) and dataset scrubbing through deduplication (Lee et al., 2022; Kandpal et al., 2022; Carlini et al., 2023a) prevent sensitive information from ever being stored in the model. However, using DP-SGD frequently results in worse generative models (Anil et al., 2022), and LMs can memorize even singleton training instances (Debenedetti et al., 2024). Scrubbing approaches are also expensive, requiring retraining for each new identified PII. To alleviate high cost incurred by retraining models, post-hoc model unlearning techniques (Liu et al., 2024) such as gradient ascent (Bourtoule et al., 2021; Neel et al., 2021; Eldan and Russinovich, 2023), reinforcement learning through human feedback (RLHF; Ouyang et al., 2022), and model editing (Cao et al., 2021; Meng et al., 2022, 2023; Wu et al., 2023) were devised. Most similar to ours is the work of Wu et al. (2023), who identify and zero out neurons responsible for memorized sensitive information. REVS diverges in key aspects, significantly outperforming naive zeroing by modifying selected neurons to demote relevant tokens in the vocabulary space, thus removing the tendency to generate sensitive information while preserving model integrity (see Section E.2).
@@ Appendix A Results
As shown in Tables 2 and 4, the Perturbation Attack (PA) score is relatively low before unlearning, suggesting that the target tokens are already highly ranked in the model’s output given the perturbed prompt. After unlearning, the PA score increases significantly. Notably, for REVS, the PA score doubles post-unlearning, indicating its effectiveness against this attack. Interestingly, on Llama-3-8B for the URL dataset, some methods (e.g., RMU and NPO) exhibit a lower PA score than the unedited model, suggesting that internal memorization persists and becomes more pronounced when the original prompt is used for unlearning with these methods.
@@ A.2 Results on Llama-3-8B
We provide the complete table of results for extraction resistance on Llama-3-8B along with the standard deviations, complementing the visual summary presented in the main paper in Table 2.
@@ A.3 Results on GPT-J-6B
REVS (ours) $99.12{\pm 3.56}$ $99.12{\pm 0.51}$ $98.55{\pm 0.2}$ $98.97{\pm 1.46}$
Emails Unedited $0{\pm 0.0}$ $0{\pm 0.0}$ $83.8{\pm 0.67}$ $44.2{\pm 4.11}$
@@
REVS outperforms all baselines on both the Unlearning Score and the Resistance Score for the GPT-J-6B model across both the SSN and Emails datasets (Tables 3 and 4). Across individual metrics, REVS consistently achieves the best or second-best performance.
@@
As shown in Figures 5, LABEL: and 4, REVS achieves superior Resistance Score, consistently outperforming all baselines including MEMIT and Head-Projection across both datasets. On the SSN dataset, REVS significantly outperforms all other baselines with near-perfect scores across all attacks. On the Emails dataset, REVS achieves best or second-best performance with comparable scores on individual attacks.
@@ A.4 Analysis
As shown in Figures 6 and 6, REVS maintains superior performance compared to MEMIT across varying candidate token set sizes ($|C|$).
@@
Figures 6 and 6 illustrate the trade-off between efficacy and specificity. REVS can achieve near-perfect specificity at low efficacy levels across both datasets, while MEMIT’s specificity saturates around $0.65$ for the Emails dataset. However, MEMIT shows more consistent specificity scores across configurations, while REVS exhibits larger efficacy-specificity trade-offs.
@@
As shown in Figures 6 and 6, both methods show declining Unlearning Scores as edit counts increase. REVS achieves near-perfect scores with few edits on the Emails dataset, while MEMIT saturates earlier. On the SSN dataset, both methods show high variance with few targets, but MEMIT demonstrates better stability with increasing targets. Higher scores are achievable by optimizing hyperparameters for larger edit counts, as current parameters were optimized for 100 targets.
Figure 5: Results for Extraction Resistance on GPTJ-6B for $k=100$. REVS is more robust to extraction attacks. Refer to Table 4 for the detailed results.
@@ B.1 Implementation and Experimental Details
For GPT-J-6B model and SSN and Email datasets, the best configuration found through this tuning process was as follows: A maximum of $N_{max}=30$ neurons were edited per layer. The neurons were chosen by first selecting the top $k=100$ neurons with the highest activation values, and then from within these neurons, the ones with the highest rank for the target token when projected to the vocabulary logit space were selected. For the target token, the desired rank $r_{h}$ within the residual hidden layer was set to $r_{h}=2600$ with an error margin of $\epsilon_{r_{h}}=1400$. The desired rank $r_{n}$ within the edited neurons was set to $r_{n}=37500$, with an error margin $\epsilon_{r_{n}}=7500$, where the maximum possible rank for any token in the GPT-J-6B vocabulary is $50400$. In our hyperparameter sweep, we considered the following ranges: the number of neurons from 10 to 1000 and the activation filter size from 10 to 2000. Additionally, we explored alternative strategies for token selection (Section E.2) and neuron selection (Section D.2).
@@ B.2 Editing Neuron
The EditNeuron step, as detailed in Algorithm 2, aims to adjust the logit value of the target token $t$ such that it achieves the desired rank $r_{n}$ in the vector of logits $\vec{v}$. This is accomplished through an iterative process that adjusts the logit value for $t$ and projects the updated logits back and forth between the vocabulary space and the hidden state space until the target rank is achieved.
@@
The process begins by initializing the logit value of $t$ to $-10$, corresponding to a low probability, and therefore low rank in $\vec{v}$. At each iteration, the logit value is adjusted by scaling it with either an increasing factor $\delta_{h}$ or a decreasing factor $\delta_{l}$, depending on whether the current rank of $t$ is lower or higher than the desired rank $r_{n}$. These factors control the adjustment rate to ensure efficient convergence. After each adjustment, the updated logits are projected back to the hidden state space and then to the vocabulary space to reevaluate the rank of $t$. This process continues until the rank of $t$ is within an acceptable margin of error $\epsilon$ relative to $r_{n}$.
@@
3:     $U^{\dagger}\leftarrow\text{Pseudo-inverse of }U$
4:     $l_{t}\leftarrow-10$triangleright Initial low logit value for target token
@@
6:         $\vec{v}\leftarrow U\vec{n}$triangleright Project neuron to logit space
7:         $\vec{v}_{t}\leftarrow l_{t}$triangleright Update logit value of $t$ in $\vec{v}$
8:         $\vec{n}\leftarrow U^{\dagger}\vec{v}$triangleright Project updated logits to neuron
9:         $\vec{v}\leftarrow U\vec{n}$triangleright Project updated neuron to logits
10:         $r_{t}\leftarrow\mathrm{Rank}(t,\vec{v})$triangleright Get rank of $t$ in projected logits
@@
14:              $l_{t}\leftarrow l_{t}*1.3$triangleright Decrease $l_{t}$ if rank is too high
@@
16:              $l_{t}\leftarrow l_{t}*0.8$triangleright Increase $l_{t}$ if rank is too low
@@
19:     return $\vec{n}$triangleright Return updated neuron value
@@ B.3 Qualitative Unlearning Examples
We demonstrate the Unlearning Effectiveness and Model Integrity of REVS through qualitative examples in Tables 5 and 6. Table 5 presents randomly sampled instances from our dataset, showcasing how the unlearning process handles private email addresses. After applying our method, the model generates plausible alternatives to the original private email addresses. Specifically, we observe two primary transformation cases: (1) replacing the original email address with a similar but different email address that maintains the contextual integrity of the original text; (2) replacing the specific email with a generic [email protected] placeholder while preserving the overall coherence and meaning of the surrounding text.
To complement our demonstration, Table 6 shows select public email addresses, exemplifies how the model maintains generation capabilities for publicly available contact information after the unlearning process.
@@ C.1 MEMIT Baseline
We used a modified version of MEMIT (Meng et al., 2023), originally proposed for inserting new knowledge into language models by editing the $FF_{2}$ matrices in model MLPs. In our case, we optimized the objective to decrease the probability of generating the target tokens instead of increasing it, to facilitate unlearning. The original MEMIT method was proposed for editing memorized facts of the form (subject s, relation r, object o), e.g., (s = Michael Jordan, r = plays sport, o = basketball), where the subject is used for calculating the updated weights. However, the targeted sensitive information might not have an explicit subject (e.g., "Send me an email with your decision soon as you get this; david.lewis@email.com"), which is the case for many samples in the Email dataset (Table 11). Furthermore, the targeted sensitive information may be memorized by various inherently different prompts where each sentence might have unique subject (e.g., "David’s email is: david.lewis@email.com" and "Contact Lewis at: david.lewis@email.com). To address this challenge, the following tuple was used as the input to the modified model: (s = last characters $n_{max\_len}$ of the prompt, r = <empty>, o = target token to unlearn).
For the Llama-3-8B model, we applied the method with specific hyperparameters: on the SSN dataset, we focused on layer $3$ with a loss break of $1\times 10^{-5}$, loss prediction probability coefficient of $100$, learning rate of $0.01$, and $5$ gradient steps with a maximum prompt length of $n_{max\_len}=20$. For the EMAIL dataset, we maintained similar settings, adjusting only the maximum prompt length to $n_{max\_len}=100$. For the GPT-J-6B model, the following hyperparameters were used: on the SSN dataset, we targeted layers $3$-$8$ with a loss break of $0.1$, loss prediction probability coefficient of $20$, learning rate of $0.05$, $25$ gradient steps, and a maximum prompt length of $n_{max\_len}=200$. The EMAIL dataset configuration was similar, with a reduced maximum prompt length of $n_{max\_len}=50$. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ C.2 FT-L Baseline
Constrained Fine-tuning (FT-L) (Zhu et al., 2020) fine-tunes the $FF_{2}$ using gradient ascent to minimize the probability of generating the target tokens $t$. FT-L imposes an $L_{\infty}$ norm constraint to limit the overall impact on the model’s knowledge. For the FT-L baseline, we employed different hyperparameter configurations across models and datasets. For both models we targeted all layers. On Llama-3-8B, For the SSN dataset, we used a loss break of $8.362\times 10^{-2}$, learning rate of $1.12\times 10^{-7}$, norm constraint of $3.3\times 10^{-5}$, and $10$ gradient steps. The EMAIL dataset configuration used a loss break of $1.225\times 10^{-3}$, learning rate of $1.75\times 10^{-7}$, norm constraint of $4.17\times 10^{-5}$, and the same $10$ gradient steps. On GPT-J-6B model, for the SSN dataset, we applied a loss break of $9.98\times 10^{-2}$, learning rate of $3.78\times 10^{-7}$, norm constraint of $7.1\times 10^{-5}$, and $10$ gradient steps across layers 00-$27$. The EMAIL dataset configuration adjusted these values, using a loss break of $1.49\times 10^{-3}$, learning rate of $1.12\times 10^{-7}$, norm constraint of $3.84\times 10^{-5}$, while maintaining $10$ gradient steps. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: learning rate from $1\times 10^{-8}$ to $5\times 10^{-3}$, loss break from $1\times 10^{-3}$ to $1\times 10^{-1}$, and norm constraint from $1\times 10^{-6}$ to $1\times 10^{-2}$, and number of gradient steps from 5 to 25.
@@ C.4 RMU Baseline
RMU (Li et al., 2024) (Representation Misdirection for Unlearning) perturbs model activations of a selected layer on the target while preserving model activations on benign data. Similar to NPO-KL, we modified RMU to target a single sensitive token. On Llama-3-8B, for SSN, we used an alpha of $1$, layer ID $7$, layers $5$-$6$-$7$, learning rate of $6.819\times 10^{-4}$, $2$ epochs, and a steering coefficient of $1$. The EMAIL dataset configuration shifted to an alpha of $1000$, maintaining the same layer and learning rate specifications, but increasing the steering coefficient to $1000$. On GPT-J-6B model, for the SSN dataset, we applied an alpha of $100$, layer ID $7$, layers $5$-$6$-$7$, learning rate of $5.19\times 10^{-4}$, $2$ epochs, and a steering coefficient of $1000$. The EMAIL dataset configuration used an alpha of $329$, the same layer specifications, a learning rate of $7.74\times 10^{-4}$, $2$ epochs, and a reduced steering coefficient of $1.61\times 10^{-1}$. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: learning rate from $1\times 10^{-5}$ to $5\times 10^{-2}$, number of epochs from 1 to 5, and values for regularization parameters $\alpha$ and steering coefficient of $[1,10,100,1000]$.

### v6

@@ 6 Related Work
Privacy concerns have given way to development of methods preventing regurgitation of PII. Differential privacy (DP; Dwork et al., 2014; Ramaswamy et al., 2020; Hoory et al., 2021; Li et al., 2022) and dataset scrubbing through deduplication (Lee et al., 2022; Kandpal et al., 2022; Carlini et al., 2023a) prevent sensitive information from ever being stored in the model. However, using DP-SGD frequently results in worse generative models (Anil et al., 2022), and LMs can memorize even singleton training instances (Debenedetti et al., 2024). Scrubbing approaches are also expensive, requiring retraining for each new identified PII. To alleviate high cost incurred by retraining models, post-hoc model unlearning techniques (Liu et al., 2024) such as gradient ascent (Bourtoule et al., 2021; Neel et al., 2021; Eldan and Russinovich, 2023), reinforcement learning through human feedback (RLHF; Ouyang et al., 2022), and model editing (Cao et al., 2021; Meng et al., 2022, 2023; Wu et al., 2023) were devised. Most similar to ours is the work of Wu et al. (2023), who identify and zero out neurons responsible for memorized sensitive information. REVS diverges in key aspects, significantly outperforming naive zeroing by modifying selected neurons to demote relevant tokens in the vocabulary space, thus removing the tendency to generate sensitive information while preserving model integrity (see Section˜E.2).
@@ Appendix A Results
As shown in Tables˜2 and 4, the Perturbation Attack (PA) score is relatively low before unlearning, suggesting that the target tokens are already highly ranked in the model’s output given the perturbed prompt. After unlearning, the PA score increases significantly. Notably, for REVS, the PA score doubles post-unlearning, indicating its effectiveness against this attack. Interestingly, on Llama-3-8B for the URL dataset, some methods (e.g., RMU and NPO) exhibit a lower PA score than the unedited model, suggesting that internal memorization persists and becomes more pronounced when the original prompt is used for unlearning with these methods.
@@ A.2 Results on Llama-3-8B
We provide the complete table of results for extraction resistance on Llama-3-8B along with the standard deviations, complementing the visual summary presented in the main paper in Table˜2.
@@ A.3 Results on GPT-J-6B
REVS (ours) $9.12{\pm 3.56} 99$$9.12{\pm 0.51}$ $98.55{\pm 0.2}$ $98.97{\pm 1.46}$
Emails Unedited ${\pm 0.0} 0±0$$\pm 0.0$} $83.8{\pm 0.67}$ $44.2{\pm 4.11}$
@@
REVS outperforms all baselines on both the Unlearning Score and the Resistance Score for the GPT-J-6B model across both the SSN and Emails datasets (Tables˜3 and 4). Across individual metrics, REVS consistently achieves the best or second-best performance.
@@
As shown in Figures˜5, LABEL: and 4, REVS achieves superior Resistance Score, consistently outperforming all baselines including MEMIT and Head-Projection across both datasets. On the SSN dataset, REVS significantly outperforms all other baselines with near-perfect scores across all attacks. On the Emails dataset, REVS achieves best or second-best performance with comparable scores on individual attacks.
@@ A.4 Analysis
As shown in Figures˜6 and 6, REVS maintains superior performance compared to MEMIT across varying candidate token set sizes ($|C|$).
@@
Figures˜6 and 6 illustrate the trade-off between efficacy and specificity. REVS can achieve near-perfect specificity at low efficacy levels across both datasets, while MEMIT’s specificity saturates around $0.65$ for the Emails dataset. However, MEMIT shows more consistent specificity scores across configurations, while REVS exhibits larger efficacy-specificity trade-offs.
@@
As shown in Figures˜6 and 6, both methods show declining Unlearning Scores as edit counts increase. REVS achieves near-perfect scores with few edits on the Emails dataset, while MEMIT saturates earlier. On the SSN dataset, both methods show high variance with few targets, but MEMIT demonstrates better stability with increasing targets. Higher scores are achievable by optimizing hyperparameters for larger edit counts, as current parameters were optimized for 100 targets.
Figure 5: Results for Extraction Resistance on GPTJ-6B for $k=100$. REVS is more robust to extraction attacks. Refer to Table˜4 for the detailed results.
@@ B.1 Implementation and Experimental Details
For GPT-J-6B model and SSN and Email datasets, the best configuration found through this tuning process was as follows: A maximum of $N_{max}=30$ neurons were edited per layer. The neurons were chosen by first selecting the top $k=100$ neurons with the highest activation values, and then from within these neurons, the ones with the highest rank for the target token when projected to the vocabulary logit space were selected. For the target token, the desired rank $r_{h}$ within the residual hidden layer was set to $r_{h}=2600$ with an error margin of $\epsilon_{r_{h}}=1400$. The desired rank $r_{n}$ within the edited neurons was set to $r_{n}=37500$, with an error margin $\epsilon_{r_{n}}=7500$, where the maximum possible rank for any token in the GPT-J-6B vocabulary is $50400$. In our hyperparameter sweep, we considered the following ranges: the number of neurons from 10 to 1000 and the activation filter size from 10 to 2000. Additionally, we explored alternative strategies for token selection (Section˜E.2) and neuron selection (Section˜D.2).
@@ B.2 Editing Neuron
The EditNeuron step, as detailed in Algorithm˜2, aims to adjust the logit value of the target token $t$ such that it achieves the desired rank $r_{n}$ in the vector of logits $\vec{v}$. This is accomplished through an iterative process that adjusts the logit value for $t$ and projects the updated logits back and forth between the vocabulary space and the hidden state space until the target rank is achieved.
@@
The process begins by initializing the logit value of $t$ to 10-10, corresponding to a low probability, and therefore low rank in $\vec{v}$. At each iteration, the logit value is adjusted by scaling it with either an increasing factor $\delta_{h}$ or a decreasing factor $\delta_{l}$, depending on whether the current rank of $t$ is lower or higher than the desired rank $r_{n}$. These factors control the adjustment rate to ensure efficient convergence. After each adjustment, the updated logits are projected back to the hidden state space and then to the vocabulary space to reevaluate the rank of $t$. This process continues until the rank of $t$ is within an acceptable margin of error $\epsilon$ relative to $r_{n}$.
@@
3:  UPseudo-inverse of $U^{\dagger}\leftarrow\text${Pseudo-inverse of }U
4:  $l_{t}\leftarrow-10$ $\triangleright$ Initial low logit value for target token
@@
6:   $\vec{v}\leftarrow U\vec{n}$ $\triangleright$ Project neuron to logit space
7:   $\vec{v}_{t}\leftarrow l_{t}$ $\triangleright$ Update logit value of $t$ in $\vec{v}$
8:   $\vec{n}\leftarrow U^{\dagger}\vec{v}$ $\triangleright$ Project updated logits to neuron
9:   $\vec{v}\leftarrow U\vec{n}$ $\triangleright$ Project updated neuron to logits
10:   $r_{t}\leftarrow\mathrm{Rank}(t,\vec{v})$ $\triangleright$ Get rank of $t$ in projected logits
@@
14:     $l_{t}\leftarrow l_{t}*1.3$ $\triangleright$ Decrease $l_{t}$ if rank is too high
@@
16:     $l_{t}\leftarrow l_{t}*0.8$ $\triangleright$ Increase $l_{t}$ if rank is too low
@@
19:  return $\vec{n}$ $\triangleright$ Return updated neuron value
@@ B.3 Qualitative Unlearning Examples
We demonstrate the Unlearning Effectiveness and Model Integrity of REVS through qualitative examples in Tables˜5 and 6. Table˜5 presents randomly sampled instances from our dataset, showcasing how the unlearning process handles private email addresses. After applying our method, the model generates plausible alternatives to the original private email addresses. Specifically, we observe two primary transformation cases: (1) replacing the original email address with a similar but different email address that maintains the contextual integrity of the original text; (2) replacing the specific email with a generic [email protected] placeholder while preserving the overall coherence and meaning of the surrounding text.
To complement our demonstration, Table˜6 shows select public email addresses, exemplifies how the model maintains generation capabilities for publicly available contact information after the unlearning process.
@@ C.1 MEMIT Baseline
We used a modified version of MEMIT (Meng et al., 2023), originally proposed for inserting new knowledge into language models by editing the $FF_{2}$ matrices in model MLPs. In our case, we optimized the objective to decrease the probability of generating the target tokens instead of increasing it, to facilitate unlearning. The original MEMIT method was proposed for editing memorized facts of the form (subject s, relation r, object o), e.g., (s = Michael Jordan, r = plays sport, o = basketball), where the subject is used for calculating the updated weights. However, the targeted sensitive information might not have an explicit subject (e.g., "Send me an email with your decision soon as you get this; david.lewis@email.com"), which is the case for many samples in the Email dataset (Table˜11). Furthermore, the targeted sensitive information may be memorized by various inherently different prompts where each sentence might have unique subject (e.g., "David’s email is: david.lewis@email.com" and "Contact Lewis at: david.lewis@email.com). To address this challenge, the following tuple was used as the input to the modified model: (s = last characters $n_{max\_len}$ of the prompt, r = <empty>, o = target token to unlearn).
For the Llama-3-8B model, we applied the method with specific hyperparameters: on the SSN dataset, we focused on layer $3$ with a loss break of $1\times 10^{-5}$, loss prediction probability coefficient of $100$, learning rate of $0.01$, and $5$ gradient steps with a maximum prompt length of $n_{max\_len}=20$. For the EMAIL dataset, we maintained similar settings, adjusting only the maximum prompt length to $n_{max\_len}=100$. For the GPT-J-6B model, the following hyperparameters were used: on the SSN dataset, we targeted layers 33-88 with a loss break of $0.1$, loss prediction probability coefficient of $20$, learning rate of $0.05$, $25$ gradient steps, and a maximum prompt length of $n_{max\_len}=200$. The EMAIL dataset configuration was similar, with a reduced maximum prompt length of $n_{max\_len}=50$. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ C.2 FT-L Baseline
Constrained Fine-tuning (FT-L) (Zhu et al., 2020) fine-tunes the $FF_{2}$ using gradient ascent to minimize the probability of generating the target tokens $t$. FT-L imposes an $L_{\infty}$ norm constraint to limit the overall impact on the model’s knowledge. For the FT-L baseline, we employed different hyperparameter configurations across models and datasets. For both models we targeted all layers. On Llama-3-8B, For the SSN dataset, we used a loss break of $8.362\times 10^{-2}$, learning rate of $1.12\times 10^{-7}$, norm constraint of $3.3\times 10^{-5}$, and $10$ gradient steps. The EMAIL dataset configuration used a loss break of $1.225\times 10^{-3}$, learning rate of $1.75\times 10^{-7}$, norm constraint of $4.17\times 10^{-5}$, and the same $10$ gradient steps. On GPT-J-6B model, for the SSN dataset, we applied a loss break of $9.98\times 10^{-2}$, learning rate of $3.78\times 10^{-7}$, norm constraint of $7.1\times 10^{-5}$, and $10$ gradient steps across layers 0-$27$. The EMAIL dataset configuration adjusted these values, using a loss break of $1.49\times 10^{-3}$, learning rate of $1.12\times 10^{-7}$, norm constraint of $3.84\times 10^{-5}$, while maintaining $10$ gradient steps. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: learning rate from $1\times 10^{-8}$ to $5\times 10^{-3}$, loss break from $1\times 10^{-3}$ to $1\times 10^{-1}$, and norm constraint from $1\times 10^{-6}$ to $1\times 10^{-2}$, and number of gradient steps from 5 to 25.
@@ C.4 RMU Baseline
RMU (Li et al., 2024) (Representation Misdirection for Unlearning) perturbs model activations of a selected layer on the target while preserving model activations on benign data. Similar to NPO-KL, we modified RMU to target a single sensitive token. On Llama-3-8B, for SSN, we used an alpha of $1$, layer ID $7$, layers 55-66-77, learning rate of $6.819\times 10^{-4}$, $2$ epochs, and a steering coefficient of $1$. The EMAIL dataset configuration shifted to an alpha of $1000$, maintaining the same layer and learning rate specifications, but increasing the steering coefficient to $1000$. On GPT-J-6B model, for the SSN dataset, we applied an alpha of $100$, layer ID $7$, layers 55-66-77, learning rate of $5.19\times 10^{-4}$, $2$ epochs, and a steering coefficient of $1000$. The EMAIL dataset configuration used an alpha of $329$, the same layer specifications, a learning rate of $7.74\times 10^{-4}$, $2$ epochs, and a reduced steering coefficient of $1.61\times 10^{-1}$. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: learning rate from $1\times 10^{-5}$ to $5\times 10^{-2}$, number of epochs from 1 to 5, and values for regularization parameters $\alpha$ and steering coefficient of $[1,10,100,1000]$.